OpenStudy (anonymous):

what is the equation of the line in slope-intercept form: 4x – 9y = –36

OpenStudy (anonymous):

slope-intercept form of a line: y=mx+c where, m=slope and c=y intercept to find the slope of a equation like ax+yb=s ----> m=-a/b and you can find the y intecept(c) by putting x=0 in this equation as y intercept is the point on the line on the y-axis where the x coordinate is always 0. now put the value of m and c in the slope intercept form to get your equation

OpenStudy (shamil98):

\[\huge 4x - 9y = -36\] \[\huge \text{To turn it to slope-intercept form:}\] \[\huge y = mx + b\] \[\huge \text{Solve for the y-variable. =)}\]

OpenStudy (balala):

Try to isolate the y on one side of the equation Hint: subtract the x variable on both sides of the equation first

OpenStudy (balala):

So we would have: -9y = -4x - 36

OpenStudy (shamil98):

\[\huge 4x - 9y = -36\] Just divide everything by -9 ( to get the y variable by itself) Then move the x to the other side to get the other form: \[\huge \frac{ -4 }{ 9 }x + y = 4\] \[\huge y = \frac{ 4 }{ 9 }x + 4\]

OpenStudy (balala):

TA DA!! :D